Circular No. 6929

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 K. Hurley, Space Sciences Laboratory, on behalf of the Ulysses Gamma-

Ray Burst Team; C. Kouveliotou, Universities Space Research Association, on behalf of the Gamma-Ray Observatory BATSE team; and E. Mazets and T. Cline, on behalf of the KONUS-WIND team, report: “This soft γ-ray repeater has entered an extraordinary new phase of activity with a previously unobserved frequency and intensity of bursts. Activity began on May 26 and has continued at least through May 30. Triangulation with Ulysses and BATSE gives a preliminary annulus whose center is at α = 21^{h}58^{m}18^{s}, δ = −11^{o}18′.4 (equinox 2000.0), and whose radius is 47^{o}.269 ± 0^{o}.019. This annulus includes the network synthesis error box (Hurley et al. 1994, Ap.J. 431, L31), the ROSAT source proposed as a possible quiescent x-ray counterpart to the soft repeater (Hurley et al. 1996, Ap.J. 463, L13), and the peculiar double M star system coincident with the ROSAT source (Vrba et al. 1996, Ap.J. 468, 225). As all these locations are mutually consistent, we conclude that the ROSAT source and M star system are indeed the counterparts to the soft γ-ray repeater, and urge that observations at other wavelengths be undertaken immediately, while the source remains active. An image may

be found at http://ssl.berkeley.edu/ipn3/sgr1900+14/.”

C. Kouveliotou, Universities Space Research Association; P. Woods, M. Kippen, and M. S. Briggs, University of Alabama in Huntsville; and K. Hurley, Space Sciences Laboratory, write: “On May 30, BATSE triggered on an extremely intense outburst from this source that comprised at least 38 bursts of varying intensity and duration; the series of bursts started at May 30.37759 UT and lasted for ∼ 350 s. The peak flux of each of the three major bursts in the series is 1.1 × 10^{−4} erg cm^{−2} s^{−1} (between 25 and 300 keV, integrated for 128 ms). Assuming a distance of ∼ 12 kpc for SGR 1900+14 (Vrba et al. 1996, Ap.J. 468, 225), this corresponds to a luminosity of ∼ 2 × 10^{42} ergs/s for each burst. This bunching of emission is very similar to the behavior of SGR 1806–20, as detected with the Rossi X-Ray Timing Explorer in 1996 (IAUC 6503). Overall, between May 26 and June 1, we have detected over 50 bursts from SGR 1900+14. The frequency and intensity of emissions are unlike any previously seen behavior from this source. We strongly encourage follow-up observations at infrared and radio wavelengths, while the source is still active.”

1998 June 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green